FIFO Method, Unit Cost, Valuation of Goods Transferred Out and Ending Work in Process Dama Company produces womens blouses a1. Prepare a physical flow schedule for April. Dama Company Physical Flow Schedule Units to account for: Units, beginning worNUMBER #3. PLEASE
Answer #1

Solution for #3

Cost of Ending Work in process $              112,000
Cost of goods Transferred out $              607,800

Working

STEP 1 (FIFO)
Reconciliation of Units
Units % already completed % completed this period
Material Conversion Material Conversion
Beginning WIP             14,000 100% 50% 0% 50%
Units introduced             65,000
Total units to be accounted for             79,000
Completed and Transferred unit             59,000 0% 0% 100% 100%
Ending WIP             20,000 0% 0% 100% 25%

.

STEP 2 (FIFO)
Equivalent Units
Total Units Material Conversion
Units Transferred: % completed this period Equivalent Units % completed this period Equivalent Units
From WIP             14,000 0%                             -   50%                          7,000
From units started/Introduced             45,000 100%                   45,000 100%                        45,000
Total             59,000                   45,000                        52,000
Total Units Material Conversion
Ending WIP % completed this period Equivalent Units % completed this period Equivalent Units
Total             20,000 100%                   20,000 25%                          5,000
Total EUP             79,000                   65,000                        57,000

.

STEP 3
Cost per Equivalent Units AND Cost to be accounted for
Material Conversion TOTAL
Cost incurred in Current Period $ 260,000 $ 364,800 $ 624,800
Total Equivalent Units                                              65,000                                                       57,000
Cost per Equivalent Units $ 4.00 $ 6.40 $ 10.40
Cost of Beginning WIP $ 25,000 $ 70,000 $ 95,000
Total cost to be accounted for $ 719,800

.

STEP 4
Cost report [Including Reconciliation]
Cost of Units Transferred               59,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Units Transferred TOTAL
A From Beginning WIP:
Material                            -   $ 4.00 $ 0
Conversion                     7,000 $ 6.40 $ 44,800 $ 44,800
B From units started & completed
Material                  45,000 $ 4.00 $ 180,000
Conversion                  45,000 $ 6.40 $ 288,000 $ 468,000
C=A+B Total $ 512,800
D Cost of Beginning WIP $ 95,000
E=C+D Total Cost of Units transferred $           607,800.00
Cost of Ending WIP               20,000.00 units
Equivalent Units [Step 2] Cost per Equivalent Units [Step 3] Cost of Ending WIP TOTAL
F Material                  20,000 $ 4.00 $ 80,000
G Conversion                     5,000 $ 6.40 $ 32,000
H = F + G Total Cost of Ending WIP $ 112000.0
I = E + H Total Cost accounted for $ 719,800
J Rounding Off Differences $ 0
K = I + J Total cost to be accounted for [equals to computed in STEP 3] $ 719,800